OpenStudy (linn99123):

HEyy buds i need help again if ya can ill medal and fan :D What is the next 3 terms in this number chain ??? 7,14,21,28,35,_,_,_

random231 (random231):

7*2=14 7*5=35 7*6= 7*7=

OpenStudy (anonymous):

it is actually just multiples of 7! 7(1) 7(2) 7(3) 7(4) 7(5) 7(6) 7(7) 7(8)

OpenStudy (anonymous):

Their correct ^^. add 7 each time.